OpenStudy (anonymous):

If the slope of the line passing through (4, -3k) and (k, 8) is 7, then find the value of k

OpenStudy (anonymous):

Help me please

OpenStudy (campbell_st):

do you know the slope formula...?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

use the slope formula then slope of the line

OpenStudy (campbell_st):

ok... it its \[m =\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] you know m = 7 so substitute the ordered pairs.... and then just solve for k.... when you have substituted the 1st step is to multiply both sides of the equation by the denominator

OpenStudy (anonymous):

Okay thank you I have another math question which is If f(x) = 2x ( 3 - 1/2x), then does f(6) = 6

OpenStudy (anonymous):

Wait * f(6) = 12

OpenStudy (campbell_st):

ummm seems a different question...?

OpenStudy (campbell_st):

I would have thought \[7 = \frac{8 + 3k}{k - 4}\] multiply both sides of the equation by the deniminator

OpenStudy (anonymous):

Yeah it's a different question